OpenStudy (anonymous):

Two friends are painting a living room. Ken can paint it in 6 hours working alone. If Barbie works alone, it will take her 8 hours. How long will it take working together? Round to the nearest tenth of an hour. A. 3.0 hours B. 3.4 hours C. 4.0 hours D. 4.6 hours

OpenStudy (anonymous):

The combined rate is the sum of the two rates. In addition, the rate is the reciprocal of the time it will take to do the job, so\[\frac{1}{6}+\frac{1}{8}=\frac{1}{t}\implies t=24/7\]Approximately 3.4 hours.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No sweat. Do you understand the algebra to get from the equation to the answer?

OpenStudy (anonymous):

yeah i wrote down the formula

OpenStudy (anonymous):

Common denominator of the LHS is 24; after adding, take the reciprocal of both sides.

OpenStudy (anonymous):

okay :D